OpenStudy (anonymous):

The radius R of a sphere is 5.4cm. Calculate the sphere's surface area A. Use the value 3.14 for "pie" and round your answer to the nearest tenth

OpenStudy (anonymous):

method is the same you need to substitute r=5.4 in, SA=4*3.14*r^2=4*3.14*5.4*5.4=366.3m^2

OpenStudy (anonymous):

check the work carefully i thing one of the digits is wrong

OpenStudy (anonymous):

is it the last one then make it 2

OpenStudy (matthewrlee):

I don't know what you're talking about but $3.14 is a great value for a pie. I say buy the pie.